Circular No. 6108

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994ae IN NGC 3370 T. Iijima, E. Cappellaro, and M. Turatto, Asiago Astrophysical Ob-

servatory, report the following observations of SN 1994ae obtained at the 1.2-m telescope of the University of Padova: “The preliminary reduction of a CCD spectrum (500–800 nm) obtained on Nov. 23.16 UT shows the lines of Si II (615 nm) and S II (529, 545 nm), indicating that the object is a type-Ia supernova near maximum light. The expansion velocity measured from the minimum of the Si II line is 11 000 km/s.”

PERIODIC COMET MACHHOLZ 2 (1994o)

Orbital elements for component A from MPC 24216:

Epoch = 1994 Sept. 5.0 TT

T = 1994 Sept.18.8041 TT  ω = 149.^{o}2623
e = 0.750145  Ω = 246.1772 }2000.0
q = 0.752548 AU  i = 12.7865

a = 3.011944 AU n^{o} = 0.1885532 P = 5.227 years

1994/95 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 24 11^{h}07.^{m}18 − 6^{o}53.^{′}4 1.200 1.288 71.^{o}4 46.^{o}6 15.7
29 11 12.43 − 8 20.6 1.212 1.344  74.6 45.0 16.3
Dec. 4 11 16.84 − 9 41.9 1.219 1.400  78.0 43.5 16.8
9 11 20.37 −10 57.4 1.224 1.456  81.7 42.0 17.3
14 11 22.98 −12 07.0 1.225 1.511  85.6 40.5 17.8
19 11 24.65 −13 10.5 1.224 1.567  89.7 38.9 18.3
24 11 25.32 −14 07.6 1.221 1.621  94.0 37.2 18.7
29 11 24.96 −14 57.9 1.217 1.675  98.6 35.5 19.1
Jan. 3 11 23.54 −15 40.9 1.212 1.729 103.5 33.6 19.6

Total visual magnitude estimates by C. S. Morris from southern California (0.26-m reflector): Oct. 7.51 UT, 9.6; 12.50, 10.2; 16.51, 11.0.

PERIODIC COMET BORRELLY (1994l) Total visual magnitude estimate by A. Hale from Descanso, CA (10×50

binoculars): Nov. 14.52 UT, 8.6.

1994 November 26 Daniel W. E. Green